### v1

# UGround: Towards Unified Visual Grounding with Unrolled Transformers

Abstract
We present UGround, a Unified visual Grounding paradigm that dynamically selects intermediate layers across Unrolled transformers as “mask as prompt”, diverging from the prevailing pipeline that leverages the fixed last hidden layer as “<SEG> as prompt”. UGround addresses two primary challenges posed by the prevailing paradigm: (1) its reliance on the fixed last hidden layer, which sequentially amplifies cumulative errors arising from layer-by-layer propagation without intermediate correction, and (2) its use of <SEG> as a prompt, which implicitly projects textual embeddings into visual space without explicit spatial cues (e.g., coordinates). Central to UGround is Policy-Prompted Masking, which comprises two key components: Stochastic Skip Connection (SSC) and Mask as Prompt (MasP). SSC is a reinforcement learning policy that, via stochastic sampling, allows each <SEG> token to slide across unrolled transformer layers, enabling dynamic layer selection at which it connects to the vision model (e.g., SAM) in a skip-connection fashion. Given the selected hidden layer, MasP uses the similarity map derived from the <SEG> token and image tokens as a soft logit mask to prompt SAM for mask generation, offering explicit spatial cues through its activation regions. To validate the effectiveness of UGround, we, for the first time, have unified visual grounding within a single framework from an attribute perspective, spanning from traditional refer expression segmentation to newly proposed reasoning segmentation, single-target to multi-target, positive query to false premise (empty target). All codes and models are publicly available at https://github.com/rui-qian/UGround.
1 Introduction
Visual grounding aims to align referring expressions with their corresponding regions in an image or video  (Mao et al., 2016; Yu et al., 2016). Despite existing efforts  (Zou et al., 2023a; Liu et al., 2024b; Zou et al., 2023b; Liang et al., 2023; Liu et al., 2023; Wang et al., 2022), it remains largely unexplored under the newly emerging task settings  (Lai et al., 2024; Wu et al., 2024; Xia et al., 2024; Rasheed et al., 2024; Ren et al., 2024; Qian et al., 2025). As shown in Table 1, the shifts across recent tasks settings reveal how visual grounding has evolved, from referring to objects by an explicit mention (RES (Rasheed et al., 2024; Ren et al., 2024)) to reasoning over descriptive language in an implicit fashion (RS (Lai et al., 2024; Yang et al., 2023)); from processing single-target queries per instance to handling multi-target scenarios (gRES (Xia et al., 2024), Multi-RS (Ren et al., 2024)); and from responding solely to positive queries to rejecting false premises with empty targets (FP-RES (Wu et al., 2024; Xia et al., 2024; Qian et al., 2025)). Such shifts reflect the intrinsic properties of tasks, which we term “attribute variation”.
However, we observe that little research has looked into unifying existing tasks within a single framework from an attribute perspective, diverging from prior works (Ren et al., 2024; Rasheed et al., 2024; Wei et al., 2025) that primarily focus on versatile capabilities. Generally, perception systems engaging with real-world scenarios are expected to reason over implicit instructions (cognition), interact with one or multiple targets (generality), and appropriately reject queries when necessary (safety). Built upon attribute basis described above, versatility-level extensions, such as grounding in videos (Wei et al., 2025) or conversations (Rasheed et al., 2024), find their footing. This leads us to ask: Can we design a unified architecture to bridge this gap?
Figure 1: Prior works typically use <SEG> token embeddings from the last hidden layer as the prompt in (a). In contrast, we leverage the similarity map in Eq. (6), generated from <SEG> and image token embeddings across dynamically selected transformer layers, as the prompt in (b). Dynamic layer selection allows the similarity map to slide across transformer layers in (c). Due to the sequential nature of transformers, $\ell_{L}-\ell_{k}$ layers are skipped, enabling a direct connection with SAM in a single forward step $\mathcal{T}_{t}$ (“skip-connection-like”). Across multiple forward steps $\mathcal{T}_{1}$ to $\mathcal{T}_{T}$, the connectivity to SAM varies dynamically, with only one path activated at each step (“dropout-like”).
Table 1: Attribute variation w.r.t. visual grounding. RES: Referring Expression Segmentation, RS: Reasoning Segmentation, FP: False Premise, gRES: Generalized RES, Multi-RS: Multi-target RS.
Method RES RS FP-RES gRES Multi-RS
LISA (Lai et al., 2024)
SESAME (Wu et al., 2024)
READ (Qian et al., 2025)
GLaMM (Rasheed et al., 2024)
GSVA (Xia et al., 2024)
PixelLM (Ren et al., 2024)
UGround (Ours)
Bearing this in mind, we revisit the methods of visual grounding presented in Table 1, where we identify two key challenges of the off-the-shelf models: (1) its reliance on the fixed last hidden layer, and (2) its use of <SEG> as a prompt. Current Large Multimodal Models (LMMs) based visual grounding typically leverage <SEG> token stemming from the fixed last hidden layer of the vision-language model’s stacked transformers (e.g., LLaVA) to prompt the downstream vision model (e.g., SAM) as shown in Fig. 1(a). These stacked transformers can be deep, e.g., 32 layers for LLava 7B, 40 layers for LLava 13B. In this paradigm, only the last hidden layer (layer 40) interacts with the vision SAM directly, whereas the other layers (layers 1-39) are separated. Consider telephone game by analogy, a message is passed from person to person (layer to layer) and becomes increasingly distorted before reaching the final participant (SAM). With only a single path, it sequentially amplifies cumulative errors arising from layer-by-layer propagation without any intermediate correction. In contrast to the similarity map with activation regions in Fig. 1(b), <SEG> token, a text placeholder, lacks explicit spatial cues on its own, e.g., coordinates. It works via the alignment from text vocabulary into vision space in an implicit way of a fully connected layer. Inspired by the above findings, a natural idea is whether we can “cheat” this telephone game by letting the final participant (SAM) tap into intermediate layers (e.g., 1–39) in advance, through “mask as prompt”.
To this end, we present UGround, a Unified visual Grounding paradigm that dynamically selects intermediate layers across Unrolled transformers as “mask as prompt”. Central to our UGround is Policy-Prompted Masking (PPM), which consists of two key components: Stochastic Skip Connections (SSC) and Mask as Prompt (MasP). As shown in Fig. 1(c), SSC is a stochastic reinforcement learning policy that allows each <SEG> token to slide across unrolled transformer layers, reinforcing dynamic layer selection at which it connects to the vision model (e.g., SAM) in a skip-connection fashion. Remarkably, SSC enables intermediate layers to skip subsequent ones and connect directly with SAM within each forward $\mathcal{T}_{t}$ (“skip-connection-like”). Across forward $\mathcal{T}_{1}$ to $\mathcal{T}_{T}$, such mechanism virtually connects all layers to SAM while activating only one pass at a time (“dropout-like”). Given the selected hidden layer, MasP uses the similarity map derived from the <SEG> token and image tokens as a soft logit mask to prompt SAM for mask generation, offering explicit spatial cues through its activation regions in the forward. Particularly, the similarity map is differentiable, beyond its implicit optimization via SAM backpropagation, we further impose a cross-entropy loss against the ground-truth mask to explicitly guide spatial cues in the backward. In summary, our contributions are threefold:
  • We have “unified” visual grounding within a single framework from an attribute perspective. Previously, existing works either focus on explicit rather than implicit expressions, or handle single-target instead of multi-target scenarios, or respond solely to positive queries without rejecting false premises (empty target).
  • We have “unrolled” stacked transformers, letting vision decoder tap into intermediate layers in a stochastic skip-connection fashion, in the way of “mask as prompt”. Importantly, we explicitly supervise the prompted mask (similarity map) against the ground truth mask to further guide the model on where to attend.
  • We conduct extensive experiments on the recently proposed ReasonSeg dataset, the classic RefCOCO(+/g) dataset, and gRefCOCO(+/g). Our UGround outperforms the state of the art, achieving up to 9.0% gains in cIoU on ReasonSeg and 12.1% gains in N-acc on gRefCOCO(+/g) val sets.
2 Related Work
Large Multimodal Models. Depending on the trade-offs between computational efficiency and performance ceiling, we categorize LMMs into two groups: (1) Fusion-centric models, which aim for deeply fine-grained multimodal interaction by projecting visual, textual, and other modalities into a shared high-dimensional space with cross-attention mechanisms (Alayrac et al., 2022; Wang et al., 2023; Pi et al., 2023; Lv et al., 2023; Zhang et al., 2024a). Flamingo (Alayrac et al., 2022) unlocks few-shot visual reasoning by bridging a frozen language model and a vision encoder with gated cross-attention. VisionLLM (Wang et al., 2023) conditions a language model on image features via cross-attention to handle diverse vision-centric tasks, generating task-specific tokens for open-ended reasoning. GPT4RoI (Zhang et al., 2024a) enables region-of-interest (RoI) reasoning through RoI-aware cross‑attention, allowing users to reference any image region in dialogue. Kosmos-2 (Peng et al., 2024) and DetGPT (Pi et al., 2023) equip LMMs with grounding capabilities. (2) Alignment-centric models, which employ a divide-and-conquer strategy, projecting the outputs of frozen or lightly fine-tuned unimodal encoders (e.g., ViT for vision, LLaMA/GPT for text) into a shared latent space via a lightweight alignment adapter (e.g., projectors, Q-Formers) (Li et al., 2023; Ye et al., 2023; Li et al., 2025; Liu et al., 2024a; Zhu et al., 2024). Otter (Li et al., 2025) instruction-tunes a Flamingo-style vision–language model on the MIMIC-IT dataset. LLaVA (Liu et al., 2024a) and MiniGPT-4 (Zhu et al., 2024) align a pretrained vision encoder with a frozen LLM using lightweight adapters for open-ended, instruction-following visual reasoning. Unlike LLaVA and MiniGPT-4, which rely on linear projections, BLIP-2 (Li et al., 2023) pioneers the Q-Former to distill visual features into compact queries for alignment with a frozen LLM. mPLUG-OWL (Ye et al., 2023) utilizes a visual abstractor to summarize global image features into condensed semantic tokens. In contrast, our UGround builds upon LLaVA (Liu et al., 2024a) for visual autoregressive modeling.
Unified Segmentation Models. Depending on the semantic granularity of joint tasks, we divide literature into two groups: Versatility-oriented segmentation, which aims to unify multiple functional properties within a single model (Wei et al., 2025; Cheng et al., 2022; Li et al., 2024; Zhang et al., 2024b; Bai et al., 2024; Rasheed et al., 2024), as opposed to prior works dedicated exclusively to semantic (Long & Shelhamer, 2015; Ronneberger et al., 2015; Badrinarayanan et al., 2017), instance (He et al., 2017; Liu et al., 2018; Carion et al., 2020), or panoptic segmentation (Kirillov et al., 2019b; Xiong et al., 2019; Kirillov et al., 2019a). Mask2Former (Cheng et al., 2022) pioneers a unified approach for image-level semantic, instance, and panoptic segmentation. OMG-Seg (Li et al., 2024) extends Mask2Former to their video-level counterparts. HyperSeg (Wei et al., 2025) and OMG-LLAVA (Zhang et al., 2024b) distinguish themselves from previous purely visual models (Cheng et al., 2022; Li et al., 2024) by leveraging LMMs to generate interactive text. VideoLISA (Bai et al., 2024) and HyperSeg (Wei et al., 2025) inject complex reasoning capabilities into generic video segmentation models. (2) Attribute-oriented segmentation, which seeks to unify the intrinsic property variations within the task itself, spanning from explicit expressions to implicit instructions, one target to many targets, and positive queries to false premises. LISA (Lai et al., 2024) elevates RES to RS by leveraging self-reasoning capabilities of LMMs. SESAME (Wu et al., 2024) and READ (Qian et al., 2025) extend LISA by rejecting false premises. GSVA (Xia et al., 2024) and PixelLM (Ren et al., 2024) generalize RES to gRES by supporting multiple targets. Closest to our work, PixelLM (Ren et al., 2024) cannot tackle empty targets, while GSVA (Xia et al., 2024) fails to handle multi-target reasoning. Attribute-oriented unification is more fundamental, as it can be applied from a versatile perspective to any specialized task. Yet, we observe that little research has looked into unifying existing tasks within a single framework from an attribute perspective, we therefore present UGround.
3 Reflection on Visual Grounding
In this section, we first revisit visual grounding with LMMs and then analyze the underlying challenges posed by the typical paradigm presented in (Lai et al., 2024; Wu et al., 2024; Qian et al., 2025).
3.1 Revisiting
Problem Definition: Let $\mathbf{x}_{img}\in\mathbb{R}^{h\times w\times c}$ denote an input image with height $h$, width $w$, and $c$ channels, and let $\mathbf{x}_{txt}$ represent the paired textual input, which can span from an explicit mention (e.g., “twigs”) to an implicit description (e.g., “what part of the tree”). For segmentation, visual grounding aims to generate a mask that localizes the region of $\mathbf{x}_{img}$ described by the textual query $\mathbf{x}_{txt}$ as
$\begin{aligned} \Theta_{\mathrm{MLE}}=\arg\max_{\Theta}\;\mathbb{E}_{\hat{\mathbf{M}}\sim\mathcal{G}_{\theta}(\cdot\mid\mathbf{x})}\big[\log\mathcal{G}_{\theta}(\hat{\mathbf{M}}\mid\mathbf{x})\big],\end{aligned}$ (1)
[…]
To equip $\mathcal{G}_{\mathcal{\theta}}$ with reasoning capabilities of LMMs, LISA (Lai et al., 2024) augments the text vocabulary of $\mathcal{G}_{\mathcal{T}}$ with a <SEG> token. Consider an input sequence $\mathbf{x}=({x}_{1},{x}_{2},\dots,{x}_{T})$, $\mathcal{G}_{\mathcal{T}}$ consumes $\mathbf{x}$ as input with $L$ stacked transformer layers, which in turn response hidden states as
$\mathcal{H}^{(\ell)}$ $=\mathcal{G}_{\mathcal{T}}^{(\ell)}\big(\mathcal{H}^{(\ell-1)}\big),\quad\ell=1,2,\dots,L,$ (2)
where $\mathcal{H}^{(0)}$ denotes the embeddings of the input sequence $\mathbf{x}$ at the embedding layer. The hidden state at the $\ell$-th layer can be presented as $\mathcal{H}^{(\ell)}=\left\{\,\bm{h}_{1}^{(\ell)},\,\bm{h}_{2}^{(\ell)},\,\dots,\,\bm{h}_{T}^{(\ell)}|\bm{h}_{t}^{(\ell)}\in\mathbb{R}^{d}\right\}$. Let $t^{*}$ be the position of the <SEG> token in the input sequence $\mathbf{x}$, then its corresponding hidden state at the $\ell$-th layer is $\bm{h}_{t^{*}}^{(\ell)}$, which is finally fed into $\mathcal{G}_{\mathcal{V}}^{dec}$ to generate the binary segmentation mask $\hat{\mathbf{M}}$ as
$\begin{aligned} \hat{\mathbf{M}}=&\;\mathcal{G}_{\mathcal{V}}^{dec}(\mathbf{f},\bm{h}_{seg}),\quad\bm{h}_{seg}=\varphi(\bm{h}^{(\ell)}_{t^{*}})\in\mathbb{R}^{d},\end{aligned}$ (3)
where $\varphi(\cdot)$ is a multilayer perceptron (MLP) projection layer, $\mathbf{f}$ is the visual features of the input image $\mathbf{x}_{img}$ extracted by the vision backbone $\mathcal{G}_{\mathcal{V}}^{enc}$ in SAM (Kirillov et al., 2023). However, we observe that existing works (Lai et al., 2024; Wu et al., 2024; Xia et al., 2024; Rasheed et al., 2024; Ren et al., 2024; Qian et al., 2025) predominantly utilize the embedding from the fixed last hidden layer ($\ell=L$) as the <SEG> token representation $\bm{h}_{seg}$, while the impact of intermediate layers remains largely unexplored.
(a) cloU on ReasonSeg val set across layers 1-40.
(b) Gauss dice loss on ReasonSeg val set across layers 1-40.
Figure 2: (2(a)) shows that the dynamic layer selection strategy outperforms the fixed last hidden layer strategy across all middle layers. (2(b)) plots the loss of similarity maps against the soft ground-truth mask, demonstrating faster convergence at intermediate layers.
3.2 Analysis
Why Dynamic Layer Selection. To analyze the impact of intermediate layers, we train UGround on the ReasonSeg train set with two strategies: (1) relying solely on the last hidden layer ($\ell=L$), and (2) dynamically selecting from the intermediate layers ($\ell\in\{1,\dots,L-1\}$). Qualitatively, we visualize the similarity maps across layers 1–40 of $\mathcal{G}_{\mathcal{T}}$ in Fig. 6 (Appendix 5). We observe that training with the fixed last hidden layer yields noisy similarity maps in intermediate layers, which eventually undermine the discriminative capacity of the final abstracted representations (6(a)). In contrast, with dynamic layer selection, the middle layers can directly align with the ground-truth mask, yielding a more discriminative foreground (6(b)). Quantitatively, we plot the performance curves of the embeddings across layers 1–40, which reveal that each layer from 10–40 outperforms the fixed last hidden layer strategy (2(a)). We also plot the per-layer similarity map loss relative to the ground-truth mask, which indicates that dynamic layer selection accelerates training convergence (2(b)).
Table 2: Quantitative analysis of the similarity maps on the ReasonSeg test split. $\mathcal{M}$prompt indicates using the mask as a prompt for the original SAM (Kirillov et al., 2023), while $\mathcal{M}$IoU quantifies the overlap (IoU) between the similarity map and the ground-truth mask. <SEG>prompt refers to using the <SEG> token as prompt for the adapted SAM, i.e., LISA (Lai et al., 2024).
Method $\mathcal{M}$prompt $\mathcal{S}$IoU <SEG>prompt
gIoU cIoU gIoU cIoU gIoU cIoU
LISA-7B (Lai et al., 2024) 16.1 17.0 32.2 35.6 47.3 48.4
GSVA-7B (Xia et al., 2024) 15.0 14.7 33.2 37.4 42.8 43.8
SESAME-7B (Wu et al., 2024) 15.6 15.6 31.6 35.0 34.9 30.7
Why Similarity as Mask. To quantitatively assess the spatial cues encoded in the similarity maps, we use them as logit masks to prompt the original SAM (Kirillov et al., 2023) for mask generation. Table 2 reveals that $\mathcal{M}$prompt is potentially capable of generating segmentation masks solely from similarity maps ($17.0\%$ vs. $48.4\%$). To further quantify the alignment between the activated regions of the similarity map shown in Fig. 6(a) and the ground-truth mask in Fig. 3, we employ the grid search-based Intersection over Union (IoU) from READ (Qian et al., 2025) to evaluate its consistency. In Table 2, $\mathcal{M}$IoU indicates that activations in the similarity map align closely with the ground-truth mask, as evidenced by a 4.3% higher cIoU on the ReasonSeg test set ($35.0\%\ \text{vs.}\ 30.7\%$) in the $3^{\text{rd}}$ row.
Summary. Analysis of the similarity maps reveals the following: (a) Similarity maps in LMMs are highly consistent with the ground-truth mask, suggesting that they can potentially serve as prompts for generating the segmentation mask. (b) Dynamic layer selection allows the intermediate layers to directly interact with SAM, resulting in a more discriminative representation, as reflected by the activated regions in the similarity maps. Motivated by the insights from the similarity maps, we leverage it in three ways: as a logit mask to prompt SAM (prompt), as a loss constraint to guide the model on where to “attend” (constraint), and as a policy reward for dynamic layer selection (signal) (see Appendix 5 for further analysis).
4 Proposed UGround
In this section, we present UGround, which dynamically selects intermediate layers across unrolled transformers and connects them to SAM in a skip-connection fashion, where the similarity map serves as a prompt, thereby contributing to a unified visual grounding paradigm. As depicted in Fig. 3, central to UGround is Policy-Prompted Masking (PPM), consisting of two components: Stochastic Skip Connection (SSC) and Mask as Prompt (MasP). SSC is a reinforcement learning policy that stochastically allows each <SEG> token to slide across unrolled transformer layers, dynamically selecting where it connects to the vision model (e.g., SAM) via a skip connection. MasP, given the selected layer, leverages the similarity map between the <SEG> token and image tokens as a soft logit mask to prompt SAM, providing explicit spatial cues through its activation regions. As the key novelty of our work is PPM, we present it first in Sec. 4.1, then the training objectives in Sec. 4.2.
4.1 Policy-Prompted Masking
Stochastic Skip Connection. Prior works (Lai et al., 2024; Wu et al., 2024; Xia et al., 2024; Rasheed et al., 2024; Ren et al., 2024; Qian et al., 2025) empirically use the embedding from the fixed last hidden layer ($\ell=L$) as the <SEG> token representation $\bm{h}_{seg}$. In contrast, we exploit intermediate layers ($\ell\in\{1,\dots,L-1\}$) to derive more discriminative semantics in a skip-connection fashion. To this end, We formulate the problem of selecting a layer for the <SEG> token as a reinforcement learning task. Each layer selection is treated as an action $a\equiv\ell^{*}$ taken in the state $s\equiv\bm{h}_{t^{*}}$, which corresponds to the unrolled <SEG> embeddings across layers. The agent receives a reward $r$ that reflects the consistency of the similarity map with the ground-truth mask. Specifically, let $\mathcal{H}_{t^{*}}$ represent the hidden states of the <SEG> token from all layers ($\ell\in\{1,\dots,L\}$) at position $t^{*}$ of the input sequence $\mathbf{x}$ as $\mathcal{H}_{t^{*}}=\{\bm{h}_{t^{*}}^{(1)},\bm{h}_{t^{*}}^{(2)},\dots,\bm{h}_{t^{*}}^{(L)}\}\in\mathbb{R}^{L\times d}$, Let $\mathbf{w}\in\mathbb{R}^{L\times d}$ denote learnable weights. The policy distribution $\pi_{\theta}(\cdot\mid\mathcal{H}_{t^{*}})$ is defined as
$\pi_{\theta}(\ell\mid\mathcal{H}_{t^{*}})=\frac{\exp(s_{\ell})}{\sum_{j=1}^{L}\exp(s_{j})},s_{\ell}=\bm{h}_{t^{*}}^{(\ell)}\cdot\mathbf{w}_{\ell},$ (4)
Figure 3: Overview of our proposed UGround. Central to UGround is Policy-Prompted Masking (PPM), which stochastically selects layer $\ell^{*}$ among “Unrolled” transformers from a policy distribution $\pi_{\theta}(\cdot\mid\mathcal{H}_{t^{*}})$ at which it connects to the vision model (e.g., SAM) in a skip-connection fashion. Given the layer $\ell^{*}$, MasP uses the similarity map as a soft logit mask to prompt SAM for mask generation, wherein we advance visual grounding within a “Unified” paradigm from an attribute perspective.
where $s_{\ell}$ is the logit scores for each layer. $\bm{h}_{t^{*}}^{(\ell)}$ is the <SEG> embedding at layer $\ell$ and $\mathbf{w}_{\ell}$ is the corresponding learnable weight. The agent chooses a layer $\ell$ according to a categorical distribution parameterized by learnable weights $\mathbf{w}$. During training, the layer $\ell^{*}$ is sampled from this distribution to allow exploration as $\ell^{*}\sim\pi_{\theta}(\cdot\mid\mathcal{H}_{t^{*}}),$ To reduce variance, we introduce a baseline $b$, which can be an EMA-smoothed reward or a critic network prediction. The actual reward $r$ is derived from the similarity map, as described in Algorithm (1) steps 5-8. We therefore formulate the REINFORCE gradient as
$\nabla_{\theta}J(\theta)\!=\!\mathbb{E}_{\ell^{*}\sim\pi_{\theta}(\cdot\mid\mathcal{H}_{t^{*}})}\Big[(r\!-\!b)\,\nabla_{\theta}\log\pi_{\theta}(\ell^{*}\!\mid\mathcal{H}_{t^{*}})\Big],$ (5)
where the log-probability of the sampled action $a\equiv\ell^{*}$ is weighted by the advantage rbr-b.
To get the reward $r$, we first compute the similarity map $\mathcal{M}$ between the <SEG> token and image tokens at layer $\ell$ and its corresponding soft mask $\mathcal{M}_{\sigma}$. Consider a sequence of discrete image tokens $\mathbf{z}=(z_{1},z_{2},\dots,z_{k})$ in the input sequence $\mathbf{x}$, where $k$ is the number of image tokens. The image token embeddings across layers are then denoted as $\mathcal{H}_{\mathbf{z}}=\{\bm{h}_{z}^{(\ell)}\mid z\in\mathbf{z},\ \ell=1,\dots,L\}\in\mathbb{R}^{k\times L\times d}$, where $\bm{h}_{z}^{(\ell)}$ is the embedding of the $z$-th image token at layer $\ell$. Given the <SEG> token embedding $\bm{h}_{t^{*}}^{(\ell)}\in\mathbb{R}^{d}$ and the image token embeddings $\mathcal{H}_{\mathbf{z}}^{(\ell)}\in\mathbb{R}^{k\times d}$ at the selected layer $\ell$, the similarity score can be computed as
$\mathcal{S}^{(\ell)}\!=\!\mathcal{H}_{\mathbf{z}}^{(\ell)}\,{h_{t^{*}}^{(\ell)}}^{\top}\in\mathbb{R}^{k},$ (6)
where each element $\mathcal{S}_{i}^{(\ell)}=(h_{z_{i}}^{(\ell)})^{\top}h_{t^{*}}^{(\ell)}$ represents the similarity between the <SEG> token and the $i$-th selected image token at layer $\ell$. $\mathcal{S}^{(\ell)}$ are arranged in a 2D grid, and interpolated to obtain the soft logit mask $\mathcal{M}\in[0,1]^{H\times W}$. Let the ground-truth mask be $M\in\{0,1\}^{H\times W}$, which we view as a function defined over the continuous 2D plane $\mathbb{R}^{2}$. By applying Gaussian smoothing, we obtain a continuous-valued function $M_{\sigma}:\mathbb{R}^{2}\to[0,1]$, where $M_{\sigma}$ represents the soft label or heatmap. Formally, $M_{\sigma}$ can be expressed as a convolution integral as $M_{\sigma}(u,v)=\iint_{\mathbb{R}^{2}}M(x,y)\,G_{\sigma}(u-x,v-y)\,dx\,dy,$ where $G_{\sigma}(x,y)\propto\exp(-\frac{x^{2}+y^{2}}{2\sigma^{2}})$, is a 2D Gaussian kernel with standard deviation $\sigma$. The reward for REINFORCE is computed using the similarity map $\mathcal{M}$ and the soft label $M_{\sigma}$ as $r=-\big(\mathcal{L}_{bce}(\mathcal{M},M_{\sigma})+\mathcal{L}_{dice}(\mathcal{M},M_{\sigma})\big),$ where $\mathcal{L}_{bce}$ and $\mathcal{L}_{dice}$ denote the Gaussian BCE and Dice losses defined over the similarity map $\mathcal{M}$. The policy loss is then defined as
$\mathcal{L}_{\mathrm{policy}}=-(r-b_{t})\,\log\pi_{\theta}(\ell^{*}\mid\mathcal{H}_{t^{*}}).$ (7)
Algorithm (1) outlines one forward pass of the PPM. As the transformer is composed of stacked layers, selecting a layer $\ell^{*}$ effectively skips the subsequent $L$-$\ell^{*}$ layers, thereby creating a direct connection to SAM and forming a skip-connection–like paradigm that mitigates the dilution of semantic information by noise in intermediate layers. Across multiple forward passes ($1$ to $\mathcal{T}$), this mechanism virtually connects all layers to SAM, with only one path active per pass in a dropout-like fashion. The current mechanism offers the following benefits: 1) Analogous to Monte Carlo Dropout, a layer $\ell^{*}$ is randomly selected in each forward pass, and multiple $\mathcal{T}$ stochastic samples facilitate an approximate Bayesian uncertainty estimation; 2) Sampling across different paths alleviates the model’s over-reliance on any single trajectory; 3) The sampling alters the model’s connectivity with SAM, virtually forming “ensemble” subnetworks. All together, these contribute to the model’s robustness.
Mask as Prompt. Compared with <SEG> tokens, the similarity map provides more explicit spatial cues, as analyzed in Sec. 3.2. We use the similarity map as a logit mask to prompt SAM. Accordingly, we reformulate Eq. (3) as $\hat{\mathbf{M}}=\;\mathcal{G}_{\mathcal{V}}^{dec}(\mathbf{f},\bm{h}_{seg},\mathcal{M}).$ The similarity map is continuously differentiable, allowing its gradients to be backpropagated through the SAM module for implicit optimization. To further guide the model to focus on target regions, we supervise the similarity map $\mathcal{M}$ by imposing a constraint (e.g., using BEC or Dice loss) against a soft ground-truth mask $M_{\sigma}$, providing explicit guidance on where the model should “attend”.
4.2 Training Objectives
Apart from the text generation loss $\mathcal{L}_{txt}$ in $\mathcal{G}_{\mathcal{T}}$, and the segmentation mask loss $\mathcal{L}_{mask}$ in $\mathcal{G}_{\mathcal{V}}^{dec}$ (Lai et al., 2024; Wu et al., 2024), we use policy loss $\mathcal{L}_{\mathrm{policy}}$ for stochastic skip connection, and binary cross-entropy (BCE) loss and Dice loss for similarity map as
$\begin{aligned} \mathcal{L}_{\mathrm{policy}}&=-(r-b_{t})\,\log\pi_{\theta}(\ell^{*}\mid\mathcal{H}_{t^{*}}),\\ \mathcal{L}_{\mathcal{M}}=\lambda_{bce}\mathcal{L}_{bce}&(\mathcal{M},M_{\sigma})+\lambda_{dice}\mathcal{L}_{dice}(\mathcal{M},M_{\sigma}),\end{aligned}$ (8)
where $M_{\sigma}$ is the soft ground-truth target, $\lambda$ denotes the loss weight, $\mathcal{M}$ is the similarity map. The overall objective $\mathcal{L}$ summarizes those losses in Eq. (8), weighted by $\lambda_{txt}$, $\lambda_{mask}$, $\lambda_{\mathcal{M}}$, and $\lambda_{policy}$ as
$\mathcal{L}=\lambda_{txt}\mathcal{L}_{txt}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}+\lambda_{policy}\mathcal{L}_{\mathrm{policy}}.$ (9)
5 Experiment
5.1 Experimental Setting
Implementation Details.
We observe that the center scaling and cropping operation in CLIP may crop out target objects. To address this issue and to align the similarity map with SAM’s input, we adopt SAM’s transformation strategy, scaling the longer side of the image to 336 and padding the image for the CLIP input $\times336$. For results on ReasonSeg, UGround-7B-LLaVA1.5 is trained on 1 NVIDIA A100 GPU 80GB for about 2 days, UGround-13B-LLaVA1.5 for about 13 hours. The training relies solely on ReasonSeg and a referring segmentation dataset ($\sim$10k images). We use LoRA (Hu et al., 2022) for efficient fine-tuning, using lora_r=8lora\_r=8 for 7B and 64 for 13B. We use the AdamW (Loshchilov & Hutter, 2019) optimizer with an initial learning rate of 3e-4, scheduled by WarmupDecayLR with 100 warmup steps. Loss weights are set as $\lambda_{mask}=\lambda_{txt}=1.0$, $\lambda_{policy}=0.1$, $\lambda_{bce}=2.0$, and $\lambda_{dice}=4.0$.
Table 3: Benchmarking reasoning segmentation models on the ReasonSeg dataset, sorted in ascending order of cIoU on the test set. *: reproduced from official models. ft: fine-tuned on 239 samples.
Method val test
overall short query long query overall
gIoU cIoU gIoU cIoU gIoU cIoU gIoU cIoU
X-Decoder (Zou et al., 2023a) 22.6 17.9 20.4 11.6 22.2 17.5 21.7 16.3
Grounded-SAM (Liu et al., 2024b) 26.0 14.5 17.8 10.8 22.4 18.6 21.3 16.4
SEEM (Zou et al., 2024) 25.5 21.2 20.1 11.5 25.6 20.8 24.3 18.7
OVSeg (Liang et al., 2023) 28.5 18.6 18.0 15.5 28.7 22.5 26.1 20.8
GRES (Liu et al., 2023) 22.4 19.9 17.6 15.0 22.6 23.8 21.3 22.0
*SESAME (Wu et al., 2024) 40.3 41.6 28.9 26.3 37.3 31.9 34.9 30.7
LLaVA1.5-7B + OVSeg (Lai et al., 2024) 38.2 23.5 24.2 18.7 44.6 37.1 39.7 31.8
*GSVA (Xia et al., 2024) 45.6 41.5 37.9 36.5 44.3 46.0 42.8 43.8
*PixelLM (Ren et al., 2024) 49.7 49.6 39.5 38.8 49.5 45.6 47.1 44.3
LISA-7B (Lai et al., 2024) 52.9 54.0 40.6 40.6 49.4 51.0 47.3 48.4
HyperSeg-3B (Wei et al., 2025) 59.2 56.7 - - - - - -
VISA-7B (Yan et al., 2024) 52.7 57.8 - - - - - -
VideoLISA-3.8B (Bai et al., 2024) 61.4 67.1 43.8 42.7 56.9 57.7 53.8 54.4
LISA-7B-LLaVA1.5 (ft) (Lai et al., 2024) 61.3 62.9 48.3 46.3 57.9 59.7 55.6 56.9
READ-7B-LLaVA1.5 (ft) (Qian et al., 2025) 59.8 67.6 52.6 49.5 60.4 61.0 58.5 58.6
LISA++-7B-LLaVA1.5 (ft) (Yang et al., 2023) 64.2 68.1 49.6 51.1 59.3 61.7 57.0 59.5
RSVP-GPT (Lu et al., 2025) 64.7 63.1 55.4 50.4 61.9 62.5 60.3 60.0
[…]
LLaVA1.5-13B + OVSeg (Lai et al., 2024) 37.9 26.4 27.1 19.4 46.1 40.6 41.5 34.1
LISA-13B-LLaVA1.5 (Lai et al., 2024) 57.7 60.3 50.8 50.0 54.7 50.9 53.8 50.8
LISA-13B-LLaVA1.5(ft) (Lai et al., 2024) 65.0 72.9 55.4 50.6 63.2 65.3 61.3 62.2
READ-13B-LLaVA1.5 (ft) (Qian et al., 2025) - - 55.4 53.7 64.4 65.1 62.2 62.8
[…]
Table 4: Benchmarking referring segmentation models on the RefCOCO(+/g) dataset, sorted in ascending order of cIoU on the RefCOCOg val set.
Method refCOCO refCOCO+ refCOCOg
val testA testB val testA testB val(U) test(U)
MCN (Luo et al., 2020) 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
VLT (Ding et al., 2021) 67.5 70.5 65.2 56.3 61.0 50.1 55.0 57.7
CRIS (Wang et al., 2022) 70.5 73.2 66.1 62.3 68.1 53.7 59.9 60.4
LAVT (Yang et al., 2022) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
X-Decoder (Zou et al., 2023a) - - - - - - 64.6 -
ReLA (Liu et al., 2023) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
SEEM (Zou et al., 2024) - - - - - - 65.7 -
Segment Anyword (Liu et al., 2025) 55.3 47.9 66.0 55.6 47.4 67.0 58.4 60.1
VISA-7B (Yan et al., 2024) 72.4 75.5 68.1 59.8 64.8 53.1 65.5 66.4
SESAME (Wu et al., 2024) 74.7 - - 64.9 - - 66.1 -
LISA-7B (Lai et al., 2024) 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
PixelLM-7B (Ren et al., 2024) 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5
READ-7B (Qian et al., 2025) 78.1 80.2 73.2 68.4 73.7 60.4 70.1 71.4
SegLLM-7B (Wang et al., 2025) 80.2 81.5 75.4 70.3 73.0 62.5 72.6 73.6
GSVA-7B (Xia et al., 2024) 77.2 78.9 73.5 65.9 69.6 59.8 72.7 73.3
OMG-LLaVA (Zhang et al., 2024b) 78.0 80.3 74.1 69.1 73.1 63.0 72.9 72.9
GLaMM-7B (Rasheed et al., 2024) 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9
POPEN-7B (Zhu et al., 2025) 79.3 82.0 74.1 73.1 77.0 65.1 75.4 75.6
[…]
Table 5: Benchmarking generalized referring expression segmentation (GRES) models on the gRefCOCO (Liu et al., 2023) dataset, sorted in ascending order of cIoU on the gRefCOCO validation set. Values are taken from (Liu et al., 2023). N-acc.: the accuracy of correct null-target classification. ft: fine-tuned on gRefCOCO training set.
Method Validation Set Test Set A Test Set B
gIoU cIoU N-acc. gIoU cIoU N-acc. gIoU cIoU N-acc.
MattNet (Yu et al., 2018) 48.24 47.51 41.15 59.30 58.66 44.04 46.14 45.33 41.32
LTS (Jing et al., 2021) 52.70 52.30 - 62.64 61.87 - 50.42 49.96 -
VLT (Ding et al., 2021) 52.00 52.51 47.17 63.20 62.19 48.74 50.88 50.52 47.82
CRIS (Wang et al., 2022) 56.27 55.34 - 63.42 63.82 - 51.79 51.04 -
LAVT (Yang et al., 2022) 58.40 57.64 49.32 65.90 65.32 49.25 55.83 55.04 48.46
ReLA (Liu et al., 2023) 63.60 62.42 56.37 70.03 69.26 59.02 61.02 59.88 58.40
LISA-Vicuna-7B (Lai et al., 2024) 32.21 38.72 2.71 48.54 52.55 6.37 39.65 44.79 5.00
GSVA-Vicuna-7B (Xia et al., 2024) 63.32 61.70 56.45 70.11 69.23 63.50 61.34 60.26 58.42
LISA-Vicuna-7B (ft) (Lai et al., 2024) 61.63 61.76 54.67 66.27 68.50 50.01 58.84 60.63 51.91
GSVA-Vicuna-7B (ft) (Xia et al., 2024) 66.47 63.29 62.43 71.08 69.93 65.31 62.23 60.47 60.56
[…]
5.2 Comparison with the State-of-the-Art
Results on ReasonSeg Dataset. In Table 3, UGround achieves competitive performance compared with the recent state-of-the-art RSVP-GPT. On the ReasonSeg validation set, UGround-7B improves over RSVP-GPT by +1.4% gIoU and +9.0% cIoU. On the test set, UGround surpasses RSVP-GPT by +3.3% gIoU and +5.4% cIoU. At the larger scale, UGround-13B further pushes the state of the art, outperforms READ-13B by up to +2.8% gIoU and +2.7% cIoU. These results validate UGround’s effectiveness in complex reasoning-oriented segmentation.
Results on RefCOCO(+/g) Dataset. In Table 4, UGround-7B achieves overall stronger performance compared with the recent state-of-the-art POPEN-7B. On RefCOCO, UGround surpasses POPEN-7B by +1.3% (val), +1.5% (testA), and +3.6% (testB), establishing clear new records across all splits. On the RefCOCOg dataset, UGround standout on test with a score of 76.1%, marking a +0.5% improvement over POPEN-7B. These results highlight that UGround consistently outperforms POPEN-7B on most benchmarks.
Results on gRefCOCO Dataset. In Table  5, UGround-7B achieves substantially stronger performance compared with the recent state-of-the-art GSVA-7B (ft) on the gRefCOCO dataset. On the validation set, UGround improvies over GSVA-7B (ft) by +5.99% gIoU, +2.27% cIoU, and +12.10% N-acc, respectively. On Test A, UGround surpasses GSVA-7B (ft) by +3.21% gIoU, +0.94% cIoU, and +8.62% N-acc. These results demonstrate that UGround not only reflects its strength in multi-object segmentation but also excels in correctly identifying empty targets (N-acc).
5.3 Ablation Study
In this section, we conduct an ablation study to analyze the contribution of each component. We report the gIoU and cIoU performance on the val set of ReasonSeg dataset.
Table 6: Ablation study on PPM.
Exp. ID SSC <SEG>prompt $\mathcal{M}$prompt gIoU cIoU
1 19.13 15.62
2 60.30 53.86
3 64.06 67.05
4 31.76 34.08
5 52.14 52.42
6 66.13 72.07
Effect of the components of PPM. In Table 6, comparing Exp. 3, Exp. 6, we observe that dynamic layer selection yields a clear improvement (+2.07% gIoU and +5.02% cIoU ), highlighting the benefit of leveraging information across multiple layers. Moreover, the decomposition indicates that using the similarity map as a prompt ($\mathcal{M}$prompt) contributes the most (e.g., Exp. 1 vs. Exp. 2, and Exp. 4 vs. Exp. 5) within this paradigm.
Table 7: Ablation study on explicit constraints for similarity map $\mathcal{M}$.
Exp. ID BCE DICE Gaussian BCE Gaussian DICE gIoU cIoU
1 64.56 71.44
2 64.57 68.63
3 63.16 65.74
4 65.80 73.49
5 63.92 68.55
6 66.13 72.07
Effect of explicit constraints for similarity map $\mathcal{M}$. In Table 7, Exp. 1, 2, and 5 optimize the similarity map against hard labels (binary masks), while Exp. 3, 4, and 6 optimize it against soft labels (continuous masks). The results demonstrate that converting the ground-truth mask into a Gaussian heatmap smooths the object boundaries, leading to performance gains of +2.21% gIoU and +3.52% cIoU when comparing Exp. 5 and Exp. 6.
Table 8: Ablation study on reward.
Exp. ID $r_{bce}$ $r_{dice}$ gIoU cIoU
1 65.37 69.53
2 63.34 68.33
3 66.13 72.07
Effect of policy reward components. We use the negative loss of the similarity map relative to the ground-truth mask as the reward. During training, it is treated as a scalar and detached from the computation graph. In Table 8, comparing Exp. 1 and Exp. 2, we observe that using $r_{bce}$ as the reward signal outperforms using $r_{dice}$, yielding a gain of +2.03% gIoU and +1.20% cIoU.
6 Conclusion
In this work, we first visualize all unrolled Transformer layers of LMMs, represented by LLaVA, and analyze the influence of intermediate layers on downstream task decoding. Building on this analysis, we present UGround, which features a dynamic layer selection and prompting mechanism—Policy-Prompted Masking (PPM). PPM enables intermediate layers to connect to the SAM decoder in a skip-connection-like manner and adjust its connectivity in a dropout-like fashion under the “mask as prompt” paradigm, together contributing to unified visual grounding. Our future work aims to further investigate how the <SEG> token interacts with the integrated SAM module for alignment.
Appendix A Supplementary Statements
Use of Large Language Models. During the preparation of this manuscript, we used Large Language Models (LLMs) primarily to aid in writing and polishing the text. Specifically, LLMs were employed for grammar correction, spelling improvements, and sentence restructuring to enhance readability and clarity. In terms of technical content, LLMs assisted only in routine tasks such as selecting tensor indices and writing dataset DataLoader code; all algorithmic ideas, methodology, and experimental designs were developed entirely by the authors. All model-generated text was reviewed and edited by the authors, who take full responsibility for the final content of this paper.
Figure 4: The UGround code repository.
The UGround code repository. As shown in Fig. 4, the UGround project is organized with a model directory for core implementations, a scripts directory containing training and testing scripts, and a configs directory for configuration files. The main model architecture is defined in UGround.py within the UGroundForCausalLM class, which extends LlavaLlamaForCausalLM and integrates a Segment Anything Model (SAM) for mask generation. A central component of this architecture is the PolicyPromptedMasking class, detailed in PPM.py, which is instantiated and used by the UGroundForCausalLM class to dynamically select hidden state features from different network layers. The selection policy is trained using the REINFORCE algorithm with either an EMA or a critic-based baseline. All codes and models are publicly available at https://github.com/rui-qian/UGround.
Figure 5: The UGround dashboard.
The UGround dashboard. As shown in Fig. 5, we developed a web-based dashboard to showcase the capabilities of UGround. The system accepts multimodal inputs, specifically a user-uploaded image paired with a natural language instruction. In response, the interface generates two outputs: (1) a segmentation mask that highlights the image region corresponding to the instruction, and (2) a textual explanation that answers the user’s query.
Visualization across layers 1-40. Fig. 68 present a qualitative analysis of similarity map activations across layers 1–40, illustrating their alignment with the ground-truth mask. Each figure contrasts heatmaps generated by prior works using a fixed last hidden layer (a) with those from our proposed Policy-Prompted Masking method (b). Unlike the baseline approach (a), our method introduces stochastic layer selection: in each forward pass, a layer $\ell^{*}$ is randomly chosen, and multiple $\mathcal{T}$ stochastic samples result in an approximate Bayesian uncertainty estimation, akin to Monte Carlo Dropout. This sampling mechanism explores diverse pathways, alleviates over-reliance on any single trajectory, and virtually forms an ensemble of subnetworks, thereby enhancing robustness. As a result, the similarity maps in (b) show activations that are sharply concentrated on the target object in intermediate and later layers, yielding more accurate alignment with the ground-truth mask compared to the diffuse and noise-prone activations observed in (a).
(a) Fixed last hidden layer (Prior Works).
(b) Policy-Prompted Masking (Ours).
Figure 6: Similarity map across layers 1-40.
(a) Fixed last layer (Prior Works).
(b) Policy-Prompted Masking (Ours).
Figure 7: Similarity map across layers 1-40.
(a) Fixed last hidden layer (Prior Works).
(b) Policy-Prompted Masking (Ours).
Figure 8: Similarity map across layers 1-40.